OpenStudy (automaticloveletter):

Which of the following is a step that can be used to find the solution to the set of equations? 2d + 1 = 3d + 5 2d = 3d + 5 2d + 1 = 3d 2d + 5 = 3d + 1

OpenStudy (automaticloveletter):

Equations: Equation A: c = 2d + 1 Equation B: c = 3d + 5 Is it A?

OpenStudy (sdfgsdfgs):

A is correct :)

OpenStudy (automaticloveletter):

The work of a student to solve a set of equations is shown: Equation A: y = 15 − 2z Equation B: 3y = 3 − 4z Step 1: −3(y) = −3(15 − 2z) [Equation A is multiplied by −3.] 3y = 3 − 4z [Equation B] Step 2: −3y = 15 − 2z [Equation A in Step 1 is simplified.] 3y = 3 − 4z [Equation B] Step 3: 0 = 18 − 6z [Equations in Step 2 are added.] Step 4: 6z = 18 Step 5: z = 3 In which step did the student first make an error? Step 4 Step 3 Step 2 Step 1

OpenStudy (sdfgsdfgs):

wat do u think?

OpenStudy (automaticloveletter):

No clue xc 1st step?

OpenStudy (sdfgsdfgs):

read them through step by step plz....its not the 1st...

OpenStudy (automaticloveletter):

2?

OpenStudy (sdfgsdfgs):

why? whats wrong there?

OpenStudy (automaticloveletter):

I'm just guessing.

OpenStudy (sdfgsdfgs):

:( but u guess right - lucky gurl!! lol

OpenStudy (automaticloveletter):

So it's the second one?

OpenStudy (sdfgsdfgs):

Step 1: −3(y) = −3(15 − 2z) [Equation A is multiplied by −3.] Step 2: −3y = 15 − 2z [Equation A in Step 1 is simplified.] See anything wrong there??

OpenStudy (automaticloveletter):

The simplification?

OpenStudy (automaticloveletter):

@sdfgsdfgs

OpenStudy (sdfgsdfgs):

yes - something went wrong there...